# Learning Individual Reproductive Behavior from Aggregate Fertility Rates via Neural Posterior Estimation

### v1

@@ 5 Results
5.1 Cross-Validation for Parameter Identifiability
Simulation-Based Inference (SBI) typically relies on summarizing potentially high-dimensional data into lower-dimensional, sufficiently informative statistics. This can be done either through explicit, pre-defined summaries or via learned embeddings. The ability to work with a low-dimensional yet descriptive representation of the data is often essential for making parameter inference tractable, especially for complex simulators.
When working with aggregate data, as in our study, the dimensionality reduction is an inherent feature of the observation process. For instance, vital statistics or administrative registers implicitly aggregate thousands of individual fertility trajectories into relatively low-dimensional summaries like age-specific fertility rates (ASFRs). We thus inherit a pre-summarized dataset, limiting our ability to engineer or learn optimal summary statistics directly from the underlying micro-level process.
A critical question then arises: do these aggregate data carry sufficient information to uniquely identify the parameters of our detailed individual-level model? Different combinations of micro-parameters (e.g., those governing desired family size, timing of intentional reproduction, or contraceptive failure) could potentially yield very similar aggregate fertility schedules, posing identifiability challenges.
To rigorously assess the degree to which our model parameters can be recovered from aggregate data, we conduct extensive validation experiments using simulated data. These experiments evaluate parameter recovery under three increasingly informative inference setups, varying the summary statistics provided to the SNPE algorithm and the strength of prior information for key parameters:
@@ 5.1 Cross-Validation for Parameter Identifiability
    ASFRs-only: Parameters are drawn from weakly informative priors. The SNPE algorithm is conditioned using only simulated age-specific fertility rates (ASFRs) as summary statistics.
@@
    ASFRs + ASUFR: Priors remain weakly informative. The summary statistics are augmented to include both simulated ASFRs and age-specific unplanned fertility rates (ASUFRs), providing direct information on the planned versus unplanned nature of births.
@@
    ASFRs + ASUFR + informative priors on $\mu_{d}$ and $\kappa$: Summary statistics are ASFRs and ASUFRs, as in Scenario 2. However, we now incorporate plausible external information by using narrower, more informative priors for the mean desired family size ($\mu_{d}$) and the baseline contraceptive failure rate ($\kappa$). All other parameters retain their weakly informative priors.
For each scenario, our validation procedure involves repeatedly generating datasets with known parameters and then attempting to recover these parameters using our full inference pipeline. Specifically, we repeat the following steps for $I$ iterations:
  1. 1.
    Generate Ground-Truth Data: In each iteration $i$ (for $i=1,\dots,I$), we draw a distinct "true" parameter vector, denoted $\theta^{*(i)}$, from the prior distribution $p(\theta)$ defined for the given scenario. Using our individual-level simulation model, we then generate a corresponding vector of aggregate summary statistics (e.g., ASFRs and ASUFRs), denoted $x_{0}^{(i)}$, which serves as the pseudo-observed data for this iteration. The pair $(\theta^{*(i)},x_{0}^{(i)})$ constitutes a ground-truth dataset where the data-generating parameters are known.
  2. 2.
    Perform Full Posterior Inference: For each pseudo-observed dataset $x_{0}^{(i)}$, we independently run our complete Sequential Neural Posterior Estimation (SNPE) procedure. This involves training a neural density estimator over $R$ rounds (in our application, $R=8$), with $N_{sim}$ new simulations (in our application, $N_{sim}=1000$) drawn from the evolving proposal distribution in each round, to obtain an approximation of the posterior distribution $q_{\phi}(\theta\mid x_{0}^{(i)})$.
  3. 3.
    Evaluate Parameter Recovery: From each resulting posterior approximation $q_{\phi}(\theta\mid x_{0}^{(i)})$, we compute a point estimate for the parameters, typically the posterior mean, denoted $\hat{\theta}^{(i)}$. We then assess the accuracy of recovery by comparing this estimate $\hat{\theta}^{(i)}$ with the known true generating parameters $\theta^{*(i)}$.
By performing these $I$ independent inference runs, we obtain a distribution of estimation errors (e.g., $\hat{\theta}^{(i)}-\theta^{*(i)}$) for each model parameter under each of the three informational scenarios. We then summarize these error distributions using the root-mean-squared error (RMSE). Lower RMSE values indicate better parameter recovery and allow for a direct comparison of how effectively each informational setup constrains the model parameters.
Figure 1 reports the normalized RMSE for all parameters under our three scenarios. The ASFRs-only scenario shows higher RMSE across parameters, indicating that ASFRs alone provide weaker identifiability for several model components. The incorporation of information on the planned versus unplanned nature of births via ASUFRs (Scenario 2) contributes to significantly reducing these errors, particularly for parameters governing the timing of reproductive intentions ($\delta_{r}$) and the rate of contraceptive failure ($\kappa$). Finally, the use of auxiliary information to inform the priors for $\mu_{d}$ and $\kappa$ (Scenario 3) yields further improvements, highlighting how complementary data can refine parameter identifiability.
@@
Figure 1: Cross-validation of eight model parameters over 50 folds. Columns 1–3 display scatterplots of true parameter values (horizontal axis) versus posterior median estimates (vertical axis) for Scenarios 1, 2, and 3, with the dashed $45^{\circ}$ reference line indicating perfect recovery. Column 4 reports the scenario-specific root-mean-squared error.
@@ 5.2 Model Validation
The cross-validation experiments (Section 5.1) confirmed the capacity of our SBI framework to recover key micro-level parameters from aggregate ASFRs and ASUFR data. To further scrutinize the model’s performance, we conduct two additional forms of validation. The first, posterior predictive checks, assesses the model’s ability to replicate the aggregate data patterns used for inference, thus evaluating its internal consistency and goodness-of-fit. The second, and more demanding, validation stage examines the model’s capacity to predict out-of-sample individual-level characteristics of the observed cohorts, such as distributions of desired family size, age at first sexual intercourse, and birth interval durations, which were not used during the parameter estimation process.
For both validation exercises, we focus on parameter posteriors obtained under Scenario 2 (ASFRs + ASUFR with weakly informative priors). This scenario is chosen for its parsimonious data requirements (relying only on the age-specific share of unplanned births alongside total ASFRs) while having demonstrated strong parameter recovery in the cross-validation exercises.
5.2.1 Posterior Predictive Checks for Aggregate Rates
@@
To perform posterior predictive checks, we drew 5,000 samples from the joint posterior distribution of the model parameters for each of the four countries (United States, Colombia, Dominican Republic, and Peru). For each parameter sample, we simulated a full cohort of individual reproductive histories using our model and then recomputed the aggregate ASFRs and ASUFRs series.
@@
Figure 2 presents the empirically observed ASFRs and ASUFRs curves against their corresponding posterior predictive distributions. In these plots, observed total ASFRs are marked by purple open circles and observed ASUFRs by purple open triangles. The red open circles and triangles indicate the mean of the 5,000 simulated ASFRs and ASUFRs schedules, respectively, while the shaded red areas depict the 97.5% posterior predictive intervals.
@@
Across all four diverse demographic contexts, the model demonstrates a strong capacity to replicate the observed data. The posterior predictive means for both ASFRs and ASUFR closely track the empirical curves, capturing key features such as the location and height of peak fertility, the age pattern of decline, and the distinct levels and shapes of unplanned fertility. For instance, the distinct, lower peak of the ASUFRs in the US is well reproduced. For Colombia (Figure 2b), the Dominican Republic (Figure 2c), and Peru (Figure 2d), characterized by higher overall fertility, the model successfully replicates not only the higher ASFRs levels but also the specific age gradients of both total and unplanned births; in Colombia and Peru, however, there is a discernible tendency for the predictive mean to underestimate the peak of the ASFRs, though the observed data remain well within the prediction intervals. Notably, across all countries, almost all empirically observed points for both ASFRs and ASUFRs fall well within their respective 97.5% posterior predictive intervals, indicating that the model, conditioned on the inferred parameters, generates aggregate outcomes highly consistent with the observed data.
@@
These results confirm that our framework, even with the relatively minimal data inputs of Scenario 2, can consistently and accurately reproduce the population-level fertility and unplanned fertility schedules it was intended to explain. This provides strong confidence in the internal validity of the estimated parameters before assessing their ability to predict out-of-sample, individual-level outcomes.
@@
Figure 2: Posterior Predictive Checks for Age-Specific Fertility (ASFR) and Unplanned Fertility (ASUFR) Rates. Comparison of observed rates (purple open circles for ASFR, purple open triangles for ASUFR) with model-based posterior predictions for four countries: (a) United States, (b) Colombia, (c) Dominican Republic, and (d) Peru. Red open circles and triangles indicate the mean of 5,000 posterior predictive simulations for ASFR and ASUFR, respectively. Red shaded areas represent the 95% posterior predictive intervals. Parameters were estimated under Scenario 2 (ASFR + ASUFR with weakly informative priors).
5.2.2 Out-of-Sample Validation
To assess the model’s ability to generate realistic individual-level reproductive behaviors not directly used in parameter estimation, we use parameter values representative of the posterior distribution (specifically, the posterior mean obtained under Scenario 2, as described in Section 5.2) to drive new simulations. These simulations yield full synthetic reproductive histories. We then compare the distributions of key micro-level outcomes from these synthetic life courses, namely, age at first sexual activity, desired family size, and inter-birth intervals, with their empirical counterparts derived from the survey data (Figure 3). This side-by-side comparison provides a direct and demanding appraisal of the model’s realism in capturing individual-level reproductive patterns.
Overall fit.
Across the four countries and three micro-level indicators, the distributions generated from the model’s synthetic life courses closely approximate the empirical distributions observed in the survey data. This indicates that the parameter vectors inferred solely from aggregate ASFRs and ASUFRs schedules enable the reconstruction of distinct micro-behavioural dimensions within the observed cohorts, without any specific tuning towards these individual-level targets. Quantitatively, for instance, the Jensen–Shannon divergence (base 2) between simulated and observed distributions never exceeds 0.056 bits for any of the three distributions evaluated.
Age at first sex.
The simulated distributions reproduce the overall shape and cross-country ordering of sexual debut: the mean age is highest in the United States and Peru, slightly lower in Colombia, and lowest in the Dominican Republic, with model-based means only a few months off the survey values. Two departures stand out: the sharp heap of reports at exactly age 18 (likely a mix of social norms and recall heaping) is smoothed out by the model, and in the United States the simulated curve is displaced very slightly to the left of the observed one, suggesting debut a few months earlier than recorded. Even with these minor mismatches, the alignment is notable, considering that the entire distribution is inferred solely from aggregate fertility rates.
Desired family size.
Overall, the simulated preferences closely match the shape of the observed distributions, reproducing the main peak around two to three children and the gradual taper into larger family ideals. A clear deviation appears at exactly two children, where the empirical data show a sharp bump (a clear outcome of strong social norms) that the model’s simple two-parameter form inevitably smooths over. Beyond this focal point the fit remains close, with only a slightly thinner extreme right tail in the simulated curves.
Birth intervals.
Together, these out-of-sample checks demonstrate that, even when estimated solely from ASFRs and ASUFRs, our mechanistic model successfully reconstructs a rich array of individual reproductive behaviors.
Age at First Sex Desired Family Size Birth Interval
United States
Peru
Colombia
Dominican Rep.
Figure 3: Validation on micro-level outcomes. Purple bars denote the observed survey distributions, while red lines show the model-implied counterparts generated from posterior draws. Columns display key behavioral dimensions—age at first sex, desired family size, and birth interval—whereas rows correspond to each study cohort (United States, Peru, Colombia, Dominican Republic).
6 Discussion
This study has introduced a powerful likelihood-free Bayesian framework capable of inferring complex individual-level reproductive behaviors solely from aggregate population data. By integrating a demographically interpretable individual-level simulation model with Sequential Neural Posterior Estimation (SNPE), we demonstrated that eight core parameters (governing reproductive timing, quantum, and contraceptive efficacy) can be reliably estimated using only age-specific fertility rates and information on unplanned fertility. The most compelling validation of our approach lies in the model’s ability, using these inferred parameters, to accurately predict out-of-sample micro-level distributions for behaviors such as age at first sexual intercourse, desired family size, and birth intervals, none of which informed the original estimation. This establishes a new, statistically rigorous bridge between macro-level fertility observations and their underlying micro-behavioral drivers.
Methodologically, this work serves as a strong empirical demonstration of SNPE’s utility for parameter inference in interpretable, yet complex, social science simulation models where direct likelihood calculation is prohibitive. The successful estimation from just two aggregate series, and the critical role of ASUFRs in disambiguating behavioral parameters and enhancing identifiability, offer transferable insights for simulation-based inference across disciplines facing similar inverse problems. Substantively, our framework provides a novel toolkit for demographers, enabling the estimation of behaviorally meaningful parameters (e.g., mean desired family size, age at intentional reproduction, contraceptive failure rates) even in data-scarce contexts, thereby offering a mechanistic complement to traditional micro-data analyses.
This capability to infer interpretable behavioral parameters significantly enhances population forecasting, moving beyond traditional methods that often rely on extrapolating aggregate trends or employ macro-level parameters with limited behavioral grounding. Our approach enables the generation of ’behaviorally explicit population forecasts’ where future scenarios are constructed from hypothesized changes in these underlying, micro-level behavioral parameters. Crucially, by first distilling complex fertility dynamics into these more causally proximate behavioral components, our framework also facilitates a clearer understanding of how external covariates, such as education or economic conditions, shape reproductive outcomes. Machine learning models, for instance, may more effectively identify systematic relationships when predicting these interpretable behavioral parameters rather than attempting to directly predict highly aggregated and confounded ASFRs. This improved ability to model the determinants of specific behaviors not only leads to more robust and theoretically sound inputs for scenario-based population projections but also enriches the potential for constructing responsive ’demographic digital twins’ for policy exploration.
Despite these advances, the present framework has limitations that define important avenues for future development. The individual-level model, while behaviorally detailed, necessarily simplifies complex human realities; notably, select dispersion parameters for reproductive preferences were fixed to ensure precise inference of core behavioral drivers from aggregate data. The crucial ASUFRs are derived from survey data on birth intentions, which carry inherent measurement considerations (such as recall bias or post-hoc rationalization), even with our methodological safeguards. Furthermore, the model’s current conceptualization—predicated on individualized fertility intentions and the distinction between planned, mistimed, and unwanted births—is best suited to contemporary populations where reproduction is substantially decoupled from sexuality and where such articulable birth control decision-making is prevalent. This may limit its direct applicability to contexts with more natural fertility regimes or where reproductive agency and intentions are structured differently. A further practical constraint is that while ASFRs are nearly universally available, the ASUFR data necessary for the full inferential power demonstrated here are far less common, currently restricting the range of settings where the complete framework can be readily applied.
These limitations notwithstanding, this research opens up several exciting directions for future work, including strategies to broaden the framework’s applicability. Addressing the contextual specificity of the current model, future efforts could adapt its structure for diverse fertility regimes. For populations closer to natural fertility, the model could be simplified, focusing on a smaller set of parameters. In contexts where Western notions of "unplanned" births are less salient, research could explore identifying and incorporating alternative summary statistics that reflect local modes of fertility regulation.
Regarding the scarcity of ASUFR data, while this study underscores the urgent need for more systematic collection of information on birth intentions, several interim strategies could enhance the model’s utility. One avenue is the development of methods to generate plausible ASUFR estimates for data-sparse regions, potentially by leveraging covariate information (e.g., education, contraceptive prevalence) from data-rich settings or by "borrowing strength" from demographically similar populations. Alternatively, the framework could be employed using ASFRs alone but with significantly stronger, empirically-grounded priors on the behavioral parameters that ASUFRs primarily help identify; this would yield posterior distributions with greater uncertainty but could still offer valuable insights, particularly for sensitivity analyses under different assumptions about underlying birth intentions.
Appendix
The visualizations presented in Figures 4 to 7 present the outcomes of the Sequential Neural Posterior Estimation (SNPE) for the United States, Colombia, the Dominican Republic, and Peru cohorts respectively. These plots were generated under Scenario 2, where model parameters were inferred using both age-specific fertility rates (ASFRs) and age-specific unplanned fertility rates (ASUFRs) with weakly informative priors. The diagonal elements of each plot display the marginal posterior distributions for the eight estimated model parameters, while the off-diagonal elements show the bivariate joint posterior distributions, revealing correlations and interdependencies between parameter pairs.
Across all four cohorts, most marginal posterior distributions are unimodal and have been substantially updated from their initial weakly informative priors (see Figure 8 for a visual comparison of the posterior distributions against their respective priors). Parameters such as $\mu_{s}$ (mean age at sexual initiation), $\sigma_{s}$ (standard deviation of sexual initiation age), $\mu_{d}$ (mean desired family size), $\kappa$ (contraceptive failure probability), and the fecundability curve parameters ($\beta_{1}$ and $\beta_{2}$) generally exhibit well-defined and relatively concentrated posterior distributions. This implies that the ASFR and ASUFR data contain sufficient information to substantially constrain these posteriors, concentrating their probability mass within a relatively confined region of the parameter space.
While most posterior distributions are well defined, $\mu_{b}$ (mean desired birth spacing) and $\delta_{r}$ (mean gap between sexual initiation and intentional reproduction) exhibit more diffuse posterior distributions in certain national contexts, indicating varying degrees of constraint by the aggregate data. Specifically, $\delta_{r}$ presents a notably wide and less defined posterior for the U.S. cohort (see Figure 4). In contrast, $\mu_{b}$ tends to be the parameter showing greater uncertainty in Colombia (see Figure 5) and Peru (see Figure 7). The Dominican Republic stands out as a setting where both $\mu_{b}$ and $\delta_{r}$ appear relatively better constrained compared to the other cohorts (see Figure 6). Despite this variability, the model’s overall predictive performance, as assessed by the posterior predictive checks (Figure 2) and the out-of-sample validation of micro-level outcomes (Figure 3), remains strong.
The off-diagonal plots in Figures 47 are essential for understanding the interdependencies between parameters. Several notable patterns emerge from these bivariate distributions.
Among these, a strong and consistent negative correlation is observed between the fecundability parameters $\beta_{1}$ and $\beta_{2}$ across national contexts. This relationship is structurally expected, as these two parameters jointly define the age-specific fecundability curve. This interdependence implies that the overall shape of the fecundability curve is a well-identified feature, even if the exact individual values of $\beta_{1}$ and $\beta_{2}$ have some trade-off. Similarly, a general negative correlation is typically observed between $\mu_{s}$ (age at sexual initiation) and $\delta_{r}$ (gap to intentional reproduction) across cohorts. This inverse relationship is also expected, as these parameters jointly determine the overall timing of entry into intentional reproduction.
A discernible negative correlation is also generally present between the fecundability peak parameter $\beta_{1}$ and the contraceptive failure rate $\kappa$ in the cohorts examined. This particular interdependency, while not as directly apparent from the model’s structure as the others, likely reflects a trade-off in fitting our observed fertility aggregates. For instance, a scenario with higher intrinsic peak fecundability (larger $\beta_{1}$) could produce similar aggregate fertility outcomes to one with lower peak fecundability if compensated by a correspondingly lower contraceptive failure rate (smaller $\kappa$), particularly when trying to match the levels of unplanned births, which are influenced by contraceptive effectiveness. This suggests that the available aggregate data may not always fully disentangle the distinct contributions of these two parameters. Nonetheless, these trade-offs are relatively contained and do not significantly reduce the model’s overall predictive utility.
Figure 4: Posterior parameter distributions for the United States cohort (Scenario 2). The diagonal displays marginal posteriors for each of the eight model parameters, while off-diagonal plots show their bivariate relationships.
Figure 5: Posterior parameter distributions for the Colombian cohort (Scenario 2). The diagonal displays marginal posteriors for each of the eight model parameters, while off-diagonal plots show their bivariate relationships.
Figure 6: Posterior parameter distributions for the Dominican Republic cohort (Scenario 2). The diagonal displays marginal posteriors for each of the eight model parameters, while off-diagonal plots show their bivariate relationships.
Figure 7: Posterior parameter distributions for the Peruvian cohort (Scenario 2). The diagonal displays marginal posteriors for each of the eight model parameters, while off-diagonal plots show their bivariate relationships.

### v2

@@ 5 Results
This section presents the results from the validation experiments designed to test our framework under the three information scenarios defined above. We first assess the model’s ability to recover known parameters from simulated data using a cross-validation procedure (Section 5.1). We then apply the framework to empirical data from our four study countries, evaluating both the in-sample fit through posterior predictive checks and, most critically, the out-of-sample predictive performance against micro-level data not used for estimation.
5.1 Cross-Validation
Our validation procedure involves repeatedly generating datasets with known parameters and then attempting to recover these parameters using our inference procedure. Specifically, we repeat the following steps for $I$ iterations:
@@ 5.1 Cross-Validation
    Generate Ground-Truth Data: In each iteration $i$ (for $i=1,\dots,I$), we draw a distinct "true" parameter vector, denoted $\theta^{*(i)}$, from the prior distribution $p(\theta)$ defined for the given scenario. Using our individual-level simulation model, we then generate a corresponding vector of aggregate summary statistics, denoted $x_{0}^{(i)}$, which serves as the pseudo-observed data for this iteration. The pair $(\theta^{*(i)},x_{0}^{(i)})$ constitutes a ground-truth dataset where the data-generating parameters are known.
@@
    Perform Full Posterior Inference: For each pseudo-observed dataset $x_{0}^{(i)}$, we independently run our complete Sequential Neural Posterior Estimation procedure. This involves training a neural density estimator over $R$ rounds with $N_{sim}$ new simulations drawn from the evolving proposal distribution in each round, to obtain an approximation of the posterior distribution $q_{\phi}(\theta\mid x_{0}^{(i)})$.
@@
    Evaluate Parameter Recovery: From each resulting posterior approximation $q_{\phi}(\theta\mid x_{0}^{(i)})$, we use the posterior mean as the point estimate for the parameters, denoted $\hat{\theta}^{(i)}$. We then assess the accuracy of recovery by comparing this estimate $\hat{\theta}^{(i)}$ with the known true generating parameters $\theta^{*(i)}$.
By performing these $I$ independent inference runs, we obtain a distribution of estimation errors (e.g., $\hat{\theta}^{(i)}-\theta^{*(i)}$) for each model parameter under each of the three informational scenarios. We then summarize these error distributions using the normalized root-mean-squared error (RMSE). Lower RMSE values indicate better parameter recovery and allow for a direct comparison of how effectively each informational setup constrains the model parameters.
Figure 1 presents the results of our cross-validation analysis, performed over 25 folds for all eleven model parameters. The first three columns display scatterplots comparing the known "true" parameter values (horizontal axis) against the posterior mean estimates recovered by our model (vertical axis) for each of the three scenarios. Perfect parameter recovery would place all points directly on the dashed $45^{\circ}$ reference line. The final column summarizes the performance for each parameter by plotting the normalized root-mean-squared error (RMSE).
The results demonstrate that parameter recovery is already reasonably robust in the baseline scenario. In the scatterplots for this scenario, the estimated parameter values consistently cluster around the true values on the 45-degree line, with some parameters exhibiting tighter clustering than others. This indicates that the model can successfully disambiguate the majority of its behavioral components from the aggregate fertility schedule alone. This is a significant finding, as it was not obvious a priori that a single aggregate series could provide sufficient information to disentangle the rich set of mechanisms in our model. As expected, parameter recovery improves systematically as more information is incorporated, with the RMSE declining for nearly every parameter in Scenarios 2 and 3.
@@
Figure 1: Cross-validation over 25 folds. Columns 1–3 display scatterplots of true parameter values (horizontal axis) versus posterior mean estimates (vertical axis) for Scenarios 1, 2, and 3, with the dashed $45^{\circ}$ reference line indicating perfect recovery. Column 4 reports the scenario-specific normalized root-mean-squared error.
@@ 5.2 Model Validation
The cross-validation experiments (Section 5.1) confirmed the capacity of our SBI framework to recover key micro-level parameters from aggregate ASFRs. To further scrutinize the model’s performance, we conduct two additional forms of validation. First, posterior predictive checks, to assesses the model’s ability to replicate the aggregate data used for inference. The second validation stage examines the model’s capacity to predict out-of-sample individual-level characteristics of the observed cohorts, such as distributions of desired family size, age at first sexual intercourse, and birth interval durations, which were not used during the parameter estimation process.
We present the main validation results using the posteriors obtained under Scenario 1 (ASFRs with weak priors). We focus on this baseline scenario as it provides the most stringent test of our model’s capacity to infer individual-level mechanisms solely from aggregate fertility rates. Results for the scenarios incorporating additional information (Scenarios 2 and 3) are presented in the Appendix.
5.2.1 Posterior Predictive Checks
@@
To perform posterior predictive checks, we draw 5,000 samples from the joint posterior distribution for each of the four countries (United States, Colombia, Dominican Republic, and Peru). For each parameter sample, we simulate a full cohort of individual reproductive histories using our model and then recompute the ASFRs.
@@
Figure 2 presents observed ASFRs against their corresponding posterior predictive distributions. Observed ASFRs are represented by purple open circles. Red open circles represent the mean of the 5,000 simulated ASFRs schedules, while the shaded red areas depict the 95% posterior predictive intervals.
@@
Across all four diverse demographic contexts, the model demonstrates a strong capacity to replicate the observed data. The posterior predictive means closely track the observed ASRFs, capturing key features such as the location and height of peak fertility, and the age pattern of decline. Notably, across all countries, almost all observed points fall well within the 95% posterior predictive interval, indicating that the model, conditioned on the inferred parameters, generates aggregate outcomes highly consistent with the observed data.
@@
These results confirm that our framework, even with minimal data inputs, can consistently and accurately reproduce the population-level fertility schedules it was intended to explain.
@@
Figure 2: Posterior Predictive Checks for Age-Specific Fertility Rates. Comparison of observed rates (purple open circles) with posterior predictions for four countries: (a) United States, (b) Colombia, (c) Dominican Republic, and (d) Peru. Red open circles indicate the mean of 5,000 posterior predictive simulations. Red shaded areas represent the 95% posterior predictive intervals. Parameters were estimated under Scenario 1 (ASFRs with weakly informative priors).
5.3 Estimated Marginal Posteriors
Figure 4 displays the marginal posterior distributions for all eleven model parameters, estimated under Scenario 1 for each of the four countries. The results clearly show that the aggregate ASFRs data provide substantial information for constraining most parameters. In nearly all cases, the posterior distributions (in red) are considerably sharper and more concentrated than their corresponding wide priors (in grey), indicating a significant reduction in uncertainty after conditioning on the observed data.
The degree to which parameters are constrained varies logically across the different components of the model. Parameters related to the timing of sexual initiation ($\mu_{s},\sigma_{s}$), the age-pattern of fecundability ($\beta_{1},\beta_{2}$), desired family size ($\mu_{d},\sigma_{d}$), and contraceptive failure ($\kappa$) are all well-constrained, exhibiting sharp, unimodal posterior distributions. The posteriors also reveal plausible cross-national differences. For example, the mean desired family size ($\mu_{d}$) for the U.S. cohort is lower than in the other three countries, as expected. Similarly, the contraceptive failure probability ($\kappa$) is estimated to be lowest in the U.S., a finding that aligns with the socio-demographic context of these cohorts. In contrast, parameters governing the timing between life-course events, like the gap to intentional reproduction ($\delta_{r}$) and the spacing between births ($\mu_{b},\sigma_{b}$), show higher posterior uncertainty. Their posteriors are wider and less distinguished from their priors, suggesting that while ASFRs alone are sufficient to identify the core components of the model, pinpointing the precise timing of intended fertility is more challenging without additional data on birth intentionality. However, it is important to note that this remaining uncertainty in specific parameters does not significantly degrade the model’s overall predictive performance, as evidenced by the successful posterior predictive checks (Figure 2). Finally, the estimation for these timing-related parameters improves substantially when direct information on birth intentionality is included, as shown in the results for Scenario 3 (see Figure LABEL:fig:appendix_posterior_comparison in the Appendix).